A flat square sheet of thin aluminum foil, 40 cm on a side, carries a uniformly distributed 255 nC charge.

a) What, approximately, is the electric field 2.0 cm above the center of the sheet? (Answer should have Units N/C)

b) What, approximately, is the electric field 19 m above the center of the sheet? (Answer should have Units N/C)
Answer #1

a) Since 2cm is << 40cm, the foil can be considered a uniformly charged infinite plane.

Then E = σ/(2ε)

σ = Q / A
E = (255nC / [0.4m]2)/(2ε)

= 9x104 N/C

b) Since 19m is >> 40cm, the foil can be considered a point charge.

Then E = kQ/(r2)

E=(9*109*255nC)/([19m]2)

= 6.357 N/C